### v1

# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

Abstract
The technological development of increasingly larger quantum processors on different quantum platforms raises the problem of how to fairly compare their performance, known as quantum benchmarking of quantum processors. This is a challenge that computer scientists have already faced when comparing classical processors, leading to the development of various mathematical tools to address it, but also to the identification of the limits of this problem. In this work, we briefly review the most important aspects of both classical processor benchmarks and the metrics comprising them, providing precise definitions and analyzing the quality attributes that they should exhibit. Subsequently, we analyze the intrinsic properties that characterize the paradigm of quantum computing and hinder the naive transfer of strategies from classical benchmarking. However, we can still leverage some of the lessons learned such as the quality attributes of a good benchmark. Additionally, we review some of the most important metrics and benchmarks for quantum processors proposed in the literature, assessing what quality attributes they fulfill. Finally, we propose general guidelines for quantum benchmarking. These guidelines aim to pave the way for establishing a roadmap towards standardizing the performance evaluation of quantum devices, ultimately leading to the creation of an organization akin to the Standard Performance Evaluation Corporation (SPEC).
I Introduction
Proper performance analysis of devices is key to provide information to potentially interested users, allowing fair comparisons among different devices. Additionally, manufacturers use these insights to keep improving their systems. The measurement of performance for classical computer systems has traditionally been controversial. Historically, it has not been easy to reach an agreement in the field on seemingly simple concepts such as what metrics are suitable to assess performance, the workload that should be used for stress tests, and even what performance means. Despite this, researchers were able to continuously improve the ability of computers to solve increasingly complex problems using an intuitive approach. This early success partially contributed to the lack of rigour in performance evaluation, and the delay of its formal introduction. Fortunately, some ground rules were set by some preliminary works on benchmarking 6448963 ; HowNotToLieWithStatistics ; Jacob1995NotesOC ; CharacterizingComputerPerformanceWithASingleDigit . These rules trained researchers in data aggregation using the most appropriate measures of central tendency, as well as they depicted common pitfalls when designing test suites. Moreover, they paved the way for the introduction of standardization through Standard Performance Evaluation Corporation (SPEC) and Transaction Processing Performance Council (TPC) HowToBuildABenchmark ; Eigenmann2001nl . Eventually, authors were able to compile a solid framework for the evaluation of performance for classical computers, offering a systematic approach to benchmarking Lilja2005o ; Jain1991-xw ; Kounev2020mh ; leboudec2010performance ; Feitelson2015 .
The adoption of standardized benchmarks within a scientific community has been shown to have a positive impact on the overall scientific maturity of a discipline Sim_BenchmarkingToAdvanceResearch . Benchmarks can help to establish standards and best practices for research and development, as well as provide a basis for objective comparison between different systems. This can lead to more rigorous and reliable research practices, and ultimately advance the field as a whole. Nowadays, the quantum computing community faces similar challenges in measuring performance as those initially encountered by classical computing researchers. Even more, quantum computing is a nascent field, with numerous technology proposals and emerging issues further complicating the situation.
The lack of standardized benchmarking rules, together with the increasing competition among major industry players vying to outperform their rivals in the quantum computing race, might lead to a situation that resonates with Goodhart’s law goodhart1975problems : ”Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes”. In other words, the absence of a rigorously formulated and scrutinized approach to benchmarking in quantum computing may distort research priorities and hinder the development of truly scalable and reliable quantum processors. This bias occurs due to benchmarking models playing a crucial role on favouring what portion of the total ”behavior space” is captured SkadronChallengesInComputerArchitectureEvaluation .
This is an already known challenge in the field of quantum computing osti_1775058 ; osti_1648785 ; osti_1639885 ; osti_1594677 ; quantumComputingSummitWhitepaperAtlanta , and there are initiatives such as the P7131 Project Authorization Request (PAR) proposal from the IEEE to standarize quantum computing performance hardware and software benchmarking StandrdForQuantumComputingPerformanceIEEE , or the proposals from IBM to define key characteristics of a quantum benchmark amico2023defining . Still, again, as the Goodhart’s law warned us, and the classic article from Dongarra et al. stated BlumeKohout2019 ; 6448963 : ”Although benchmarks are essential in performance evaluation, simple-minded application of them can produce misleading results. In fact, bad benchmarking can be worse than no benchmarking at all.” Therefore, we believe there is a clear need for a standardized benchmarking framework in the field of quantum computing to be able to assess how good the different attempts are at measuring performance.
In this article, we propose general guidelines to steer practitioners in the development of fair and good quantum benchmarks and drive quantum processor development in the right direction. Our methodology focuses on incorporating some lessons learned over decades from classical computer benchmarking theory, adapting them to the intrinsic features, limitations, and potentials of quantum computing. In this sense, we analyze the fundamental elements of both classical benchmarks and the metrics that comprise them, identifying the quality attributes they must exhibit to be considered good. We evaluate the main quantum metrics and benchmarks proposed in the literature to determine which satisfy these quality attributes. Additionally, we introduce precise definitions for commonly used terms in quantum benchmarking to ensure proper use of terminology. Finally, we introduce a benchmarking structure that provides an overview of the most important factors to be considered in benchmark creation. Our efforts aim to advance towards the standardization of the evaluation of quantum device performance.
This article is organized as follows. In Sec. II, we provide an overview of the key elements of both classical processor benchmarks and metrics, emphasizing the fundamental properties they must exhibit to be considered good benchmarks and metrics. Additionally, we establish a common definition of various concepts used in the field of classical benchmarking. Finally, we review some examples of classical benchmarks and their limitations. Then, in Sec. III, we introduce some of the distinctive features of quantum computing devices that distinguish them from classical counterparts. We unify and define concepts to be used in the field of quantum benchmarking based on the definition of their classical counterpart. Furthermore, we review existing proposals for assessing the performance of a quantum computer, presenting the most relevant techniques, metrics, and benchmarks, and assessing which quality attributes these metrics and benchmarks fulfill. In Sec. IV, we propose general guidelines for any practitioner to follow when proposing a quantum benchmark. We also compile a series of open questions that need resolution on the path toward standardizing the comparison of quantum processor performance. Subsequently, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC). Finally, in Sec. V, we present our concluding remarks.
II What can we learn from Classical benchmarking?
Originally, the word benchmark used to refer to ”a mark on a workbench used to compare the lengths of pieces so as to determine whether one was longer or shorter than desired” specglossary . Even though the term has clearly evolved, it retains the intuitive idea of comparing one object against another one to measure its properties. Hence, it is not a surprise that the concept of computer benchmark  [^1] We focus in this article on the study of benchmarking related to physical devices or processors, even though interesting debates are held by the community in other benchmarking fields such as algorithms benchmarking BenchmarkingOptimization ; TestingHeuristics . Indeed, similar approaches to the ones proposed for classical algorithms benchmarking could be leveraged by the quantum computing community too, for scenarios such as assessing heuristics or the overhead of implementing quantum algorithms in real devices. is based on this very same idea, and it is usually defined as ”a test, or set of tests, designed to compare the performance of one computer system against the performance of others” specglossary .
A natural question that then arises is what we mean by performance. Performance is typically conceived in the classical computing scope as the useful work that a computer can get done in a given amount of time, and/or with a set of given resources. Indeed, in the past, it was common to hold the premise that ”the time required to perform a specified amount of computation is the ultimate measure of computer performance” CharacterizingComputerPerformanceWithASingleDigit . For classical computers, there have been historically two major approaches to the performance analysis problem: (1) Experimenting and measuring the actual system, and (2) Experimenting with a model of the system. For the latter, for any kind of performance evaluation not limited to computing devices you can use a physical model or a mathematical model. For computational performance assessment, we prefer to build models that can represent the system in terms of quantitative and logical relations. These models can later be either solved using mathematical tooling to obtain an analytical solution, or simulated Zeigler2018fg ; SimulationModelingAndAnalysisLawAverill . Even though in this work we will focus on the measurements of experiments with the actual systems, it is possible that we will see other approaches involving system modeling in the future. So far, and to the best of our knowledge, models have only been used by the quantum community for the assessment of quantum networks Wu2021 ; Coopmans2021 , still there are other event-based models of quantum phenomena published DeRaedt2005 ; Willsch2020 .
The aforementioned definitions already illustrate some of the inherent complexity of benchmarking. Besides these, the multidimensional nature of modern computational systems only exacerbates the problem. For instance, some systems may exhibit greater performance when constrained to a small set of resources, while others may perform better when given a longer total execution time. Also, the tests themselves must be appropriately designed and specified to ensure that they do not introduce performance biases. As an example, a particular test program may heavily rely on floating-point operations, while another may focus on intensive integer arithmetic, leading to performance biases that favor certain architectures over others on modern computer devices.
In order to accurately evaluate the performance of devices, it is important to consider their various characteristics. Based on past attempts, it has become evident that a comprehensive assessment requires a holistic approach that takes into account multiple factors CharacterizingComputerPerformanceWithASingleDigit . Furthermore, the tests used in such assessments must meet certain standards of fairness to ensure that the results are not biased. Therefore, a rigorous and well-defined methodology, that is specifically tailored to the device being evaluated, is required. Failure to do so could result in inaccurate or misleading conclusions.
In this section, we will provide an overview of the fundamental characteristics that a quality benchmark should exhibit. This will be followed by a discussion of the different types of benchmarks that can be developed. Additionally, we will address the properties that a good metric should exhibit for benchmarking purposes. Furthermore, we will define relevant concepts used in the field of classical benchmarking. Finally, we will examine various instances of classical benchmarks and their associated constraints.
II.1 Characteristics of a good benchmark
Before delving into the characteristics that define a good device/processor benchmark, and the importance of considering these factors when creating and using them, it is important to establish a common definition of the term. Although there are various interpretations of the term, we will agree with the definition of the term proposed by the Standard Performance Evaluation Corporation (SPEC) specglossary :
Definition 1 (Classical Benchmark).
A benchmark is a test, or set of tests, designed to compare the performance of one computer system against the performance of others.
This terminology has been widely accepted in the field and provides a comprehensive definition of what a benchmark should involve. It is worth noting that other sources have also proposed alternative definitions for the term Lilja2005o ; SkadronChallengesInComputerArchitectureEvaluation ; Kounev2020mh ; HowToBuildABenchmark ; leboudec2010performance ; Jain1991-xw . We would like to remark that a benchmark is chosen in order to test the performance with respect to a given test suite, i.e. an application or set of applications. A better performance for this task does not necessarily mean a better performance for other tasks.
According to the definition, the crux of creating a high-quality benchmark rests in selecting appropriate tests that precisely reflect the types of workloads the system is expected to encounter, or that measure the anticipated behavior of the system during specific tasks relevant to the assessor. In other words, the tests should be representative of the performance of the system in real-world scenarios (i.e. the portion of the behavior space that is considered relevant) and should provide a realistic measure of the capabilities of the system related to the task. On classical benchmarking, researchers have been deeply studying how to model the workload so that it is representative of the behavior it models. Note that this definition implies that the workload is context-agnostic, but that is not the case. A workload is only a good fit for a problem when restricted to a certain context (e.g., imagine a workload given by a sequence of operations. If the operations can be executed in parallel by the device, or if the performance metric is not correlated to the devices capacity of solving operations, the workload might not be useful) Feitelson2015 . For classical devices, special detail has been paid to understanding the main contributing factors that need to be considered when modelling a workload, and multiple examples can be found in the literature ComputerResearchWorkload ; ArtificialWorkloadDesign . Indeed, in Ref. Jain1991-xw , R. Jain explicitly states that the discussion on selecting the most appropriate workload is the deepest pitfall into which a performance analysis project may fall. Alternatively, in a simplistic manner, we can state that the workload is ultimately the input for the tests, and hence the old computer science classic garbage in, garbage out (GIGO) principle applies.
Apart from characterizing the workload, the benchmark has to provide flexibility in both its implementation and usage to accommodate different hardware configurations, software environments, and use cases. This flexibility should extend to the ability to configure and customize the tests, as well as to the option of running them on a variety of systems, platforms, and architectures. By incorporating these features, a benchmark can provide a reliable and accurate measure of system performance and enable meaningful comparisons among different systems.
To be able to build a good benchmark, we introduce below a list of quality traits to ensure that it is comprehensive, accurate, and representative of system performance. These attributes are in line with those proposed by Kounev Kounev2020mh and also reflect the most desirable characteristics of benchmarks proposed by other authors in the field HowToBuildABenchmark ; GustafsonAndSnell_HINT ; Henning_SpecCpu2000 ; Huppler_TheArtOfBuildingGoodBenchmark ; Sim_BenchmarkingToAdvanceResearch .
  1. 1.
    Relevance: A benchmark should closely correlate with behaviors that are of interest to users and provide relevant information about the aspects of performance of interest. Some examples are speed, scalability, accuracy, energy consumption, or quality.
  2. 2.
    Reproducibility: A benchmark should produce consistent results when run with the same test configuration, enabling users to replicate and verify results.
  3. 3.
    Fairness: A benchmark should allow for different test configurations to compete on their merits without introducing biases or artificial limitations.
  4. 4.
    Verifiability: A benchmark should provide confidence that its results are accurate and can be verified using well-established and transparent methodologies.
  5. 5.
    Usability: A benchmark should be easy to use and the cost of running it should be affordable in order to avoid roadblocks for users to run it.
An in-depth analysis of the different features presented above can also help us to understand how to better design benchmark models, and why they are important.
In particular, let us start by considering relevance, the first key attribute of an ideal benchmark. Crafting a benchmark that correlates with the performance of the system on a specific domain poses a considerable challenge, as it often results in limited applicability beyond that particular area. Conversely, benchmarks designed with the intention of broader applicability tend to sacrifice the depth of meaningfulness in a specific scenario. Thus, achieving the optimal balance of relevance in benchmark design requires careful deliberation and trade-offs, ensuring that the benchmarks scope aligns with the intended application domain while maintaining a meaningful level of specificity Huppler_TheArtOfBuildingGoodBenchmark .
Regarding reproducibility, we see that even though it may initially appear to be a straightforward feature to attain, it often proves to be hard to meet in practice. A system or process is deemed to be in a steady state when the variables governing its behavior remain constant over time and benchmarks typically operate in a steady state. We can model this definition by assuming that the properties of the workload/test of the benchmark are random variables $X_{i_{1}},\dots,X_{i_{n}}$ on given instants $i_{1},i_{2},\ldots,i_{n}$. Then, a system is said to be in a steady state (stationary state) whenever the joint distribution $X_{i_{1}+s},\dots,X_{i_{n}+s}$, where $s$ is a shifting index, matches the original one. By taking $n=1$ and $n=2$, it is possible to prove that the properties must have the same expected value, variance, and covariance. To model process in which the properties are independent from each other, we can relax the condition to define weak stationary states. This can be achieved by only enforcing $\mathbb{E}[X_{i}]=k$ (all properties have the same expected value), and the covariance of $X_{i}$ and $X_{j}$ for any $i,j$ is a function satisfying that $\text{Cov}(X_{i},X_{j})=\gamma(|i-j|)$. Either steady-state/stationary or weak stationary are enforced as prerequisites to achieve meaningful results in the performance analysis.
Typical applications of the devices translate into variations in load due to factors such as user usage patterns. This disparity between benchmarking and real-world applications underscores the difficulty in achieving full reproducibility. The steady-state problem impacts other factors too such as fairness, by potentially introducing fluctuating biases. In general, replicating the dynamic conditions of actual usage becomes complex. As we will see in the next sections, this aspect becomes even more pertinent when considering quantum devices that lack substantial mitigation and error correction techniques, requiring frequent recalibration. A simplistic model of this non-stationary behavior for each property of the workload could be $X_{i}=C_{i\,\text{mod}\,T}+Z_{i}$ where $Z$ is the error component, $C$ is the cyclic part (defined as $C_{t}$ for $0\leq t\leq T$), and $T$ is the period or cycle time for the calibration.
As we have introduced, fairness can be substantially influenced by the conditions before executing the benchmark. Even more, fairness requires ensuring that multiple configuration rules can be applied when running tests while ensuring a common ground is established to allow comparisons. To facilitate fair comparisons, on classical computing a mechanism known as peak and base metrics were introduced by SPEC CPU2017 cpu2017 . Base metrics require that the test programs, constituting the building modules of a test suite in a given language, are compiled using the same rules (i.e. compilation flags), in the same order. Also, these metrics are mandatory for the generation of the benchmark report. On the other hand, optional peak metrics can be included allowing for greater flexibility, enabling the usage of different compilation rules for each benchmark, as well as feedback-directed optimization.
Similarly, verifiability presents comparable challenges that we should carefully consider too. In industry, benchmarks are typically run by vendors who have a vested interest in the outcomes. In contracts, in academia results are subjected to a peer review process to be repeated and built upon by other researchers. Good benchmarks should introduce self-validation, since this may improve verifiability and usability, ensuring that results can be independently corroborated and effectively utilized.
Finally, usability is an important feature that a benchmark should satisfy. When a benchmark is easy to use, it is more likely to be adopted by users, organizations, and industry.
Even a meticulously designed benchmark incorporating the aforementioned key features is susceptible to manipulation by different agents with conflicting goals. As an illustrative example, in Ref. Jain1991-xw , R. Jain presents a thought-provoking discussion on strategies employed to gain an advantage in what he terms the ratio game. A ratio game is a technique that involves using ratios with incomparable bases and skillfully combining them to benefit oneself. This is a common scenario for benchmarking, whenever we want to assess performance among distinct devices, normalizing against the performance of a reference device. For this, even when the geometric mean is properly used to average the normalized values HowNotToLieWithStatistics , it is possible to derive rules for beating the ratio game. Upon analysis, three rules emerge regarding the ratio game:
  1. 1.
    If a system is better on some benchmarks and worse on others, contradicting conclusions can be drawn in some cases.
  2. 2.
    If one system outperforms all others across all benchmarks, the application of any ratio game technique fails to yield contradictory conclusions.
  3. 3.
    Even when one system exhibits better performance than another on all benchmarks, the selection of an appropriate base  [^2] In a ratio game, we are comparing devices performance using ratios, where the denominator is called the base. can be manipulated to showcase a more favorable relative performance.
These statements highlight the susceptibility of benchmarks to manipulation, emphasizing the need for cautious interpretation and critical evaluation when drawing conclusions based on the results from ratio games. Other researchers such as Mashey have also call out the difficulties of aggregating data, and the controversy of using one one Pythagorean mean over the others, as well as stated the common agreement that ”performance is not a single number” Mashey2004 . Nonetheless, even though there is agreement on the fact that a single number is delusional, there are still discussions to determine what is the preferred metric, sparking debates over which numerical representations are deemed more optimal in each context.
Finally, it is important to remember that while a good benchmark can provide a reliable and accurate measure of system performance, it cannot solve the problem of evaluating conflicting aspects of a computer device. Instead, it can help to properly balance the trade-offs for a particular objective and influence its strengths and weaknesses. It is not possible for a single benchmark to excel in all areas, and there is no silver bullet to solve this problem SkadronChallengesInComputerArchitectureEvaluation . Therefore, multiple workloads and benchmarks are necessary to comprehensively evaluate the performance of a system. Each benchmark may focus on different aspects of performance and may use different metrics and methodologies. At best, we can obtain a set of benchmarks that will be representative of the behavior space we want to model and measure. By combining the results of these multiple benchmarks, it would be possible to obtain a comprehensive and accurate picture of the system. Ultimately, the choice of benchmark(s) should depend on the specific use case and the objectives of the evaluation, as the relevancy requirement already stated.
II.2 What types of benchmarks are there?
Benchmarks can be classified according to multiple criterias, ranging from the objective of what portion of the behavior space they want to assess to the way the workload is defined. All of these categories fall under one of the following three different strategies when trying to assess the performance Lilja2005o ; Kounev2020mh :
  • Fixed-work workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of time required to perform a fixed amount of computational work, as seen in the case of the old classic Dhrystone and Whetstone benchmarks. These benchmarks compiled a list of programs to be executed, the former favouring floating point operations, while the latter focuses on integer performance analysis WhetstoneBook ; DhrystonePaper .
  • Fixed-time workloads: This benchmarking strategy refers to the types of benchmarks that focus on measuring the amount of computation performed within a fixed period, as seen in the case of the SLALOM benchmark. This benchmark runs an algorithm to calculate the radiosity, with the performance metric being the accuracy of the answer computed within a fixed time SLALOMBenchmark .
  • Variable-work and variable-time workloads: This benchmarking strategy refers to the types of benchmarks that allow both the amount of computation performed and the time to vary. A great example is the HINT benchmark that aimed to find rational upper and lower bounds for $\int_{0}^{1}\frac{1-x}{1+x}.\,dx$. The assessment was later performed by calculating the quality improvements per second (QUIPS), i.e. dividing the quality of the solution by the execution time GustafsonAndSnell_HINT .
In general, the performance of a system depends on the performance of multiple components of the system that are utilized during a certain operation. The problem with using fixed-work benchmarks is that they introduce an intrinsic performance bottleneck when measuring the increment of improving a single component of the system. Benchmarking models have been developed changing the workload type precisely to address this concern, and further discussions on overcoming issues with the fixed-work via using fixed-time benchmarks can be found in Ref. Gustafson1992 ; Isoefficiency . However, benchmarks that allow for variability in both time and workload provide greater flexibility to evaluate the performance of different systems, determining which machine produces the highest-quality result in the shortest time possible.
One approach to benchmarking classification is obtained by focusing on the level of implementation required from usersHowToBuildABenchmark ; Kounev2020mh :
  • Specification-based benchmarks define a problem or functions that must be implemented, the required input parameters, and the expected outcomes. The responsibility for implementing the specification falls upon the individual conducting the benchmark.
  • Kit-based benchmarks include the implementation as a necessary part of the official benchmark execution.
  • Hybrid benchmarks provide some functions as part of the kit, while others are allowed to be implemented at the discretion of the individual conducting the benchmark.
The specification-based benchmarks are sometimes called ”pen and paper” benchmarks, and they have proved successful when there are considerable differences in the architectures on which the benchmarks will be implemented ReflectionsOnCreationParallelBenchmarkSuite . Still, they come at the cost of ad-hoc and expensive implementations from their users, as well as challenges with verifiability. On the other hand, kit-based benchmarks reduce the cost and time to run them by offering execution-ready implementation, which favoured them historically. Indeed, the stone aged benchmarks for classical computing are good examples of this, such as the Whestone, Dhrystone or the Linpack benchmarks WhetstoneBook ; DhrystonePaper ; OverviewOfCommonBenchmarks .
Another approach to benchmarking classification is obtained by focusing on the workload Kounev2020mh ; Lilja2005o :
  • Synthetic benchmarks, when artificial programs are constructed to try to mimic the underlying characteristics of a given class of applications.
  • Microbenchmarks, small programs used to test specific parts of the system, independent of the rest (e.g. evaluate the I/O (input/output) subsystem in a computer).
  • Kernel benchmarks or program kernels, small programs that capture the key portion of a specific type of application (e.g. execute the portion of a program that consumes a larger fraction of the total time within an application).
  • Application benchmarks, complete real programs designed to be representative of a particular class of applications.
Finally, we must not forget that to be relevant, benchmarks must correlate with the expected behavior that wants to be assessed. Hence, there are different benchmarks that deal with different goals in classical computing. For instance, we could benchmark memories, GPU, Networks, Databases, Power Consumption, CPU-intensive operations, and each of them might require customised benchmarks to properly capture its behavior space.
II.3 What makes a good metric
As discussed, a benchmark is a test or workload that we characterise according to certain criteria, we must always later measure the results of each test. The measurement of these results can range from the time it takes to finish to more complex metrics such as the quality of a solution. To further break down how to properly define these metrics, let us start first introducing the term:
Definition 2 (Performance Metric).
A value derived from the evaluation of a function that maps a considered set of objects or events to another set of values, characterizing a given performance-related property of a system.
From Lilja and Kounev Lilja2005o ; Kounev2020mh , we know that these can be typically classified into three groups: Counting how often a certain event occurs (e.g. number of operations processed in a given observation interval), duration of an interval (e.g. time taken to complete a single operation), and size of a parameter (e.g. precision obtained to compute a value). Some examples of basis performance metrics are the response time (i.e., time taken by a system to react to a request providing a respective response), throughput (i.e., rate at which requests are processed by a system), and utilization (i.e., fraction of time in which a resource is used).
Similarly to the quality attributes defined for a benchmark, there are properties proposed in the literature for an ideal performance metric Kounev2020mh ; Lilja2005o ; HowToBuildABenchmark :
  1. 1.
    Practical. A metric is practical if it can be measured in a viable and realistic time. This is complicated to define, but we will call so if it can be efficiently computed, i.e. within polynomial time in terms of the relevant size of the problem, thus facilitating its practical applicability. We understand that for some polynomials and sizes it can be challenging, but we must add a threshold. Additionally, this will help meet the usability criteria for the benchmarks using it.
  2. 2.
    Repeatable. If the metric is measured multiple times using the same experiment, the same value must be measured. Ideally, a metric should be deterministic when measured multiple times. In practice, small differences are acceptable, provided they are bounded. Note that this will allow the benchmark to meet the reproducibility criteria.
  3. 3.
    Reliable. It must rank systems consistently with respect to a property that is subject to evaluation. In other words, if system $A$ performs better than system $B$ with respect to the property under evaluation, then the values of the metric for the two systems should consistently indicate so. The reliability of a metric is closely correlated with the relevancy of a benchmark.
  4. 4.
    Linear. Linear metrics are intuitively appealing since humans typically tend to think in linear terms: twice as high value of the metric should indicate twice as good performance.
  5. 5.
    Consistent. It has the same units and the same precise definition across different systems or configurations of the same system. When this property is not met, the benchmark will suffer poor verifiability and potentially lack of relevancy.
Not all of the historically ideal properties of a metric must be preserved for a metric to be good enough to build a good benchmark. For instance, linearity is not preserved for the Ritcher scale but this does not prevent us from leveraging it as an useful metric. On the other hand, independent metrics could act as a sufficient but not necessary condition for the lack of influence on the metric to favour certain manufacturers. In general, when a metric does not fulfill a certain condition, it is actually possible to compensate it with other metrics within the same benchmark.
We can also classify metrics depending on what they measure. Metrics that measure what was done, useful or not, have been called means-based metrics whereas end-based metrics measure what is actually accomplished. As an example, for a given program we could measure the floating point operations per second (FLOPS), or the total execution time. The former is a mean-based metric, while the latter is an end-based metric. Depending on the context, it could be argued that one is better than the other Lilja2005o .
Other well-known examples of classical metrics are clock rate, or million instructions per second (MIPS). The clock rate is repeatable (given a fixed processor), practical, consistent, and independent metric. However, given its limited scope, the clock rate is both non-linear and unreliable as a performance metric (a processor with a faster clock rate does not necessarily imply better overall performance. The clock rate ignores many important performance-relevant aspects such as how much computation is actually performed in each clock cycle as well as the interaction with the memory and I/O subsystems). On the other hand, MIPS is neither reliable, linear, or consistent (for the latter, take the example of different Instruction Set Architectures, such as RISC vs CISC IsaWars ). Ultimately, we mentioned before FLOPS. Unfortunately, this metric is neither reliable, linear, consistent, nor independent (processor manufactures may use different rules for what is counted as floating-point operations, which makes the metric fail for consistency and independence).
Finally, it is important to notice that the simple act of measuring a property using a metric might be tough. The effect of the Simpsons’ paradox serves as an adage to the designers of benchmarks and metrics, calling out the possibility of reversing or disappearing the trend on the data when combining it leboudec2010performance .
II.4 Examples of classical benchmarks
Classical computing benchmarking has evolved to deal with multiple and distinct applications, from Machine Learning processing to High Performance Computing (HPC). In each of these fields, ad-hoc efforts are required to define the proper benchmarks. As an example, the original LINPACK benchmark based on subroutines for analyzing and solving systems of simultaneous linear algebraic equations and linear least squares problems has been used as the base for the HPLinpack highly parallel computing benchmark for HPC evaluation dongarra1979linpack ; Dongarra2003TheLB . Indeed, the Top 500 list top500 ranks the most powerful non-distributed computer systems using HPLinpack. Interestingly, there are still discussions about the suitability of this approach for HPC benchmarking CriticsTop500 . For instance, the HPLinpack benchmark overrates the available compute cycles for real-world applications. Consequently, alternatives such as the High Performance Conjugate Gradients (HPCG) benchmark are being explored as an alternative for evaluating supercomputers for the TOP 500 list (unfortunately, even the alternatives such as the HPCG are not immune to techniques that artificially inflate the problem size to rig the results) PerformanceModelingHPCG .
Other examples of controversial benchmarks used historically are the synthetic ones, such as the Whetstone WhetstoneBook . Even though it was proposed to measure the floating point performance of a CPU, it can be misleading when compared with real-life situations. Nowadays, computers can easily fit the main loop of Whetstone in the low-level caches of the CPU, keeping the math coprocessor pipeline filled and achieving its optimal speed. Moreover, they face the issue of most synthetic benchmarks, they are not guaranteed to be correlated with real application scenarios, and fail to meet true relevancy.
The issues with Whetstone as well as LINPACK showed the need for updating and adapting benchmarks to keep up with the advances in technology. A possible way to deal with this is the approach taken by the benchmarks developed by Standard Performance Evaluation Corporation. These benchmarks are updated regularly. What is more, they have been developing custom-made benchmarks for differents types of applications. For instance, SPECviewperf 2020 allows users to measure graphics performance gwpg.spec.org while SPECaccel 2023 helps when assessing the performance of computationally intensive parallel applications running under the OpenACC and OpenMP target offloading APIs spec.org.accel2023 , or SPEC CPU 2017 that enables comparison of compute-intensive workloads resolution capacity on different computer systems (including optional metrics to measure energy consumption)spec.org.cpu2017 .
III Quantum benchmarking
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including quantum annealing for optimization problems; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems; and digital quantum computing, which uses a gate-based quantum devices to simulate a quantum system. These three paradigms offer different perspectives on the practical applications of quantum computing.
The diversity of technological platforms and paradigms used in the development of quantum computers introduces significant complexity in comparing the performance among different devices. For this reason, the search for a common benchmark that allows comparison of quantum computers performance across different technological platforms and paradigms poses a challenge to the scientific community. Despite these challenges, efforts have been made to develop benchmarks enabling comparison among diverse quantum computers. Primarily, these efforts focus on three types of relevant comparisons: comparing the performance of different quantum processors within the same computational paradigm (e.g., comparison of digital quantum computing platform providers such as Forest, Qiskit, ProjectQ LaRose2019overviewcomparison , or comparisons of trapped atomic ions and superconducting circuits Linke2017 ); comparing different computational paradigms (e.g., between processors for digital quantum computing and quantum annealing or digital-analog quantum computing Flannigan_2022 ); and finally, comparing quantum computers to classical computers (e.g. comparing digital processing against quantum annealing Jnger2021 , empirical comparisons relying upon simplified Machine Learning (ML) tasks havenstein2018comparisons ; Cugini2023 or even the claims on quantum supremacy Arute2019 ; kalai2022google ). It is worth noting that these three comparisons are crucial for advancing scalable and reliable quantum technologies, but each of them demands a thorough analysis. Furthermore, a similar analysis could be expanded to include algorithmic comparisons. Looking ahead, benchmarking should be extended to cover other potential new aspects of a quantum computer that we have not even envisioned yet (e.g. qRAM QuantumRam or other future components that we have not even thought about yet). During this work, we will restrict ourselves to the evaluation of the performance of quantum processors.
Despite technological advances, we are still undergoing the development of the quantum computing foundations. As Blume-Kohout and Young stated osti_1594677 , we are still more in the early stages of the quantum computing technology world rather than close to any exponential growth, which translates into challenges as we do not have clearly defined applications for these devices. What is more, we lack knowledge of how each the performance of each component will propagate and how the different quantum technology architectures will evolve. The current state is closely related to classical devices in the pre-1940 stages with the introduction of devices that were able to mechanize sequential series of calculations and procedures in binary code via the Bell Labs Model I, which served as an example of successfully applying Boolean logic to the design of computers TROPP1980115 . Back then, computers were starting to be explored, in contrast to upcoming years such as 1965 when Moore published the famous paper that led to the establishment of the Moore’s law MooreLaw , and exponential growth took off. This contrasts with the ”benchmarketing” approach that we sometimes see in the community when promoting the results of quantum computing experiments.
In the quantum computing field, we have deceptively misleadingly used the term benchmarking. While it traditionally refers to a series of tests evaluating the performance of a device, its application has expanded to encompass the metrics evaluated as part of the test, the actual tests, or even for verifying the success of a test. We believe it is crucial to establish a unified terminology for quantum benchmarking to provide a solid foundation for comprehensive and effective evaluations of quantum systems.
In this section, we introduce some of the intrinsic characteristics of quantum computer devices that make them different from classical ones, making the benchmarking model non-transferable. Additionally, we review the technological stages that we expect these devices will undergo throughout their evolution. Subsequently, we propose a unified terminology for the quantum benchmarking field. Finally, we review some existing proposals metrics and benchmarks for quantum processor in the literature, assessing what quality attributes they satisfy.
III.1 Nuances of benchmarking quantum computers
There is lack of a definite computational paradigm, standard, platform, or even architecture for quantum computers. This is an eventual consequence of our limited knowledge so far about the real scope of quantum computing devices and platforms. Note that, in classical computing, regardless of the type of architecture (i.e. Princeton/Von Neumann vs. Harvard), there is a similar unified approach that allows for performance evaluation using well-established metrics, workloads, and even benchmarks. Indeed, it can be argued that differences in the architectures have been historically deceptive mythHarvard , and that devices can be easily compared assuming a similar common ground. Interestingly, this was not always the case. As mentioned before, back in the times of the Bell Labs Model I, every computer device was custom and hard to compare against each other. Benchmarking was not so extended and generalised as a recommended best practice until the 70s-80s.
Similar to these early years, quantum computers are a field in its first stages. Based on our current expectations, we could clearly divide the future different quantum technological development eras as it follows:
  1. 1.
    Noisy Intermediate-Scale Quantum (NISQ) Preskill2018 , characterized by middle-size quantum processors with tens to hundreds of noisy physical qubits, substantially limiting the size of the quantum circuits that can be reliably executed. In the current state of NISQ era, it is controversial whether quantum processors have already reached the necessary technological development to achieve quantum advantage.
  2. 2.
    Partially Quantum Error Corrected (PQEC), when quantum processors errors are below a certain threshold that allows the application of error correction codes, such as the surface code Fowler2012 . When an insufficient number of physical qubits are available, we could employ a smaller code and partially correct errors up to a certain distance-code and number of errors per per cycle of stabilizer measurements Andersen2020 .
  3. 3.
    Fault Tolerant Quantum Computers. Once sufficiently large-distance codes may be implemented, then there will be theoretically no limit in the size of the quantum circuits and the full computational power of quantum computers will be deployed.
  4. 4.
    Fully Functional Quantum Computers. Existing quantum algorithms and circuits consider the use of ancillary qubits, which play a similar role as cache memory in classical processors. However, for Fully Functional Quantum Computers, there will also be necessary the existence of Quantum Random Access Memories (qRAM) QuantumRam , and even long-term (error-corrected) quantum memories Terhal2015 .
We live currently in the NISQ era, while the community is still investing to keep advancing quantum computers to the next stages. To meet the PQEC era, we need to further explore error correction codes. It is important to consider that the number of physical qubits required by the error corrections codes to correct any error generally depends on how below the threshold the processor and operations are Andersen2020 ; suppressingQuantumErrors . The generation of quantum processors developed in this era could be considered a transition age towards a fault-tolerant quantum computer in which, in principle, there are no limitations in the size of the quantum circuits that can be reliably executed. It is important to remark that these codes rely on sufficiently trustworthy qubits and operations, not only quantum gates, but also very accurate measurements.
The challenges that we are facing today within the NISQ era are different from the ones we will be facing in the upcoming eras, and hence the benchmarks should reflect this. Given the wide diversity of technological platforms and paradigms for implementing quantum computers, it is likely that benchmarks in the NISQ era will focus on comparing the performance among different quantum computers within subsets of the broad existing variety. As technology advances, it is expected that the wide variety of existing technological platforms will naturally decrease, preserving only those that are most competitive for the implementation of quantum computing. In this context, as technology continues to progress, comparing the performance of quantum computers implemented on different technological platforms will become more feasible. On the other hand, achieving a fair comparison of performance between quantum computers and classical computers at present is not viable, as quantum computers require different workloads, metrics, and overall approaches to benchmarking than classical computers. This is expected to occur in the era of Fully Functional Quantum Computers. In Sec. IV, we will discuss further what recommendations we suggest to properly face the different quantum technological development eras.
As mentioned at the beginning of this article, quantum computing exhibits fundamental characteristics and limitations associated with the nature of quantum physics, rather than being tied to technological development. Consequently, classical benchmarks and metrics cannot be directly extended to quantum computing. We now summarize some of these intrinsic characteristics of quantum computing:
Figure 1: Input-Processing-Output (IPO) model. IPO model can be applied to many quantum algorithms, with quantum state preparation as input, quantum information processing as data processing, and quantum state tomography as information retrieval.
  • The fundamental unit of information in a quantum computer is the qubit (quantum bit). While classical bits can only represent a single binary value 0 or 1, qubits have the ability to exist in multiple states simultaneously, thanks to a phenomenon known as superposition. As a result, a qubit can be expressed as a linear combination of both the 0 and 1 states.
  • Even though an $n$-qubit quantum state can in principle contain an exponential quantity of information in bits 10.1145/581771.581773 , the amount of classical information that can be retrieved by a projective measurement is limited to, at most, $n$ bits due to Holevo’s bound Maslov_2021 . As a consequence, if the algorithm requires to retrieve the full quantum state, i.e. to perform full tomography of the state, then an exponential number of measurements will be required.
  • After every run of the algorithm, if the information is classically extracted, the processor must be restarted in the initial state. This state is typically represented as $|\psi_{0}\rangle=\otimes_{i=1}^{n}|0\rangle_{i}$, where $n$ denotes the number of qubits. This process is usually slow and has to be taken into account in the running time of the algorithm.
  • A quantum computer execution typically sticks to the Input-Process-Output (IPO) model, a rather general framework used in software engineering to describe the structure of an information processing program. Generally, a quantum computer needs to complete three different stages: the input or loading of information, the processing thereof, and the output or retrieval of relevant information. Most quantum algorithms adhere to this model  [^3] There are notable exceptions, such as hybrid classical-quantum algorithms., where the input is state preparation or matrix exponentiation, and the output is just the quantum measurement (see Fig.1), adding extra steps to the classical algoritmic execution.
  • On a quantum computer, after processing, we have a quantum state that needs to be measured in a basis. Measurement involves projecting the quantum state using a series of projectors that collapse the state to one of the bases (single-shot measurement). To obtain the probabilities of each state, this process must be averaged over $N$ repetitions, where $N$ depends on the precision with which we aim to recover the coefficients and the number of qubits comprising the state. More specifically, as every run is independent of the previous ones, the variance in the estimation is asymptotically given by the Cramér-Rao bound $(\Delta p)^{2}\geq(FN)^{-1}$, with $F$ the Fisher information associated to the aforementioned projective measurement.
  • While classical computers programs executions can be optimised by the compilers, quantum computers can be optimised during the transpilation process of the quantum circuit (e.g. translating a Qiskit circuit into OpenQASM) as well as during the translation of the circuit into the device custom control logic to operate the device (e.g. translating OpenQASM into microwave pulses for a superconductor quantum computer). This means that there are two level of abstractions that offer potential for optimizations, and therefore can impact the performance assessment of a device.
Apart from these are fundamental features associated to quantum physics, there are other relevant aspects of quantum computing which mainly affect the performance for the NISQ and PQEC eras.
  • The number of qubits in quantum computing is a crucial factor that directly influences the capabilities and potential of quantum computers. Increased quantities of qubits empower the quantum computer to solve more complex problems. However, it is not only the quantity of qubits that is important, but also their quality.
  • Qubit connectivity plays a key role in the performance of quantum computing because it determines the interactions between the qubits. To perform quantum computations, qubits need to interact through quantum gates. When qubits are not directly connected additional operations, referred to as swap operations, must be applied to facilitate their interaction. These additional operations increase the depth and runtime of the quantum algorithm, affecting the performance of noisy quantum computers, especially in NISQ and PQEC eras.
  • Noise has a significant impact on the performance of quantum computing. Quantum computers are inherently susceptible to errors and inaccuracies due to various sources of noise such as thermal environment and external control factors. The effects of noise affect different aspects of quantum computation. High gate error rates reduce the fidelity of quantum operations, affecting the accuracy of computations. Furthermore, the interaction with an environment causes decoherence, leading to the loss of quantum information. High decoherence levels limit the circuit depth that can be accurately executed on the quantum computer, directly affecting the complexity of the algorithms that can be executed. Furthermore, NISQ devices require periodic recalibration due to the effect of the thermal environment and the external control.
The number of qubits available and their quality, the connectivity, and the noise will have a greater impact on NISQ and PECQ devices than on the later stages for the assessment of the performance. The routing and mapping of an algorithm will play a key role in the performance of a quantum computer, as simple changes on how the algorithms are executed can lead to drastically different results even for the same hardware.
III.2 A common language to discuss
The lack of a common and accepted standard for defining techniques and processes in quantum computing has led to the misuse of terminology. Even though there has been past attempts to classify the different types of benchmarks, and review approaches in the field, the terms quantum benchmarking, validation, verification, and testing have been loosely and inconsistently used in the literature. While we appreciate the groundwork laid by amico2023defining in establishing a foundation for benchmarking standards, our approach extends beyond insights solely derived from quantum computing research efforts. We aim to incorporate not only the lessons learned from the quantum domain but also draw upon the extensive findings from classical computing research. In this subsection, we will propose a unified terminology for the main concepts used in the quantum benchmarking field.
To ensure alignment on terminology, we precisely define below some of the terms commonly used in quantum benchmarking, with definitions inspired by the classical ones:
Definition 3 (Quantum Benchmark).
A test, or set of tests, that aims to measure or evaluate the performance, efficiency, or other properties of a quantum processor or hardware component for a certain task.
The underlying goal of a quantum benchmark is to be able to compare devices with each other. What is more, similarly to classical benchmarking, quantum benchmarks involve running specific tests or workloads on the systems and software collecting data to analyze its performance. QASMBench li2022qasmbench , Quantum Volume PhysRevA.100.032328 , or Quantum LINPACK QuantumLINPACK are examples of quantum benchmarks.
When executing these benchmarks, we need to later compile a list of metrics that will describe the system, we introduce the definition of a performance metric analogous to the classical case:
Definition 4 (Quantum Performance Metric).
A quantum performance metric is a performance metric, Definition 2, where the property belongs to a quantum system.
When the contextual information provides the details about the system being assessed, we briefly refer to them as just metrics. Interestingly, Quantum Volume is not only a benchmark but also a metric PhysRevA.100.032328 , even though it is tightly coupled to the Quantum Volume benchmark definition. Other examples of metrics are cross-entropy difference, $\ell_{1}$-Norm Distance to measure total divergence of probability distributions to the sample space, the gates fidelity, the number of qubits, or the Q-Score Martiel2021BenchmarkingQC .
Following is the definition of quantum verification, which refer to the process of certifying that a quantum device works as it is expected:
Definition 5 (Quantum Verification).
Process of evaluating whether or not a quantum computer is indeed producing correct results.
This concept is particularly relevant, but it comes with challenges. In particular, if the approach to verify correctness lays on the assumption that the device would be able to execute correctly the output of certain input instances, we need to be able to answer the following question Gheorghiu2018 : ”If a quantum experiment solves a problem which is proven to be intractable for classical computers, how can one verify the outcome of the experiment?”. Or, as the Aaronson’s \$25.00 prize states AaronsonPrize : ”If a quantum computer can efficiently solve a problem, can it also efficiently convince an observer that the solution is correct? More formally, does every language in the class of quantumly tractable problems (BQP) admit an interactive proof where the prover is in BQP and the verifier is in the class of classically tractable problems (BPP)?”. Unfortunately, no one has claimed the prize yet. Still, there are other approaches worth flagging in the field leveraging quantum provers as trusted measurement devices Mahadev2018 .
Related to verification, we can find the concept of verification techniques for quantum benchmarks:
Definition 6 (Quantum verification technique).
Any procedure that aims to produce efficiently verifiable results of a quantum experiment.
In this category, we can find approaches such as the circuit mirroring Proctor2021 , or quantum state tomography PhysRevA.40.2847 , which can be leveraged to help quantum benchmark designers when trying to measure the results of their computations.
We need to clarify that quantum verification and the quantum benchmarking verification techniques are different from the classic concept of formal verification of programs. Indeed, there exists a parallel research field dedicated to the formal verification of quantum programs FormalVerificationQuantumPrograms , and we can easily extend the notion from their classic counterparts. Note that even though the formal verification of programs is an interesting field with intensive research for both classical and quantum cases, historically a testing-based approach that does not rely on any mathematical formalization for the proof of correctness has been favoured. For those cases, we introduce the idea of testing quantum programs:
Definition 7 (Quantum Testing).
Process of checking the behavior of a quantum program to confirm it matches expected requirements.
There are some early works in the field of testing GarcadelaBarrera2021 . Some of them raise awareness about the challenges of introducing classical computing techniques for testing such as white- and black-box testing 8805685 , and hence the need for an enhanced approach for quantum computers. The importance of quantum testing will become greater in the upcoming years as we transition to the next eras.
Finally, we would like to call out that there could be approaches to define benchmarks that allow the definition of families of circuits or programs, or they can provide a specification for a subfamily of those. For these cases, we introduce the definition below:
Definition 8 (Quantum Benchmark Framework/Family).
A structure or specification on how to define a family of tests for a benchmark.
The most well-known example in this category would be the volumetric benchmark framework Blume_Kohout_2020 .
Also, as mentioned in this article, it is important to highlight that no single metric can capture the whole behaviour of any classical or quantum computing device. Hence, to overcome the danger of focusing too much on a single aspect, the concept of benchmark suites can help us:
Definition 9 (Quantum Benchmark Suite).
A collection of quantum benchmarks, generally aiming to model different aspects of the behavior of a quantum computer.
III.3 Reviewing existing proposals for benchmarking
Recently, there have been attempts to categorize benchmarking practices. However, the lack of a precise definition of concepts has led to confusion in the field of quantum benchmarking. There are instances where the concepts of certification and benchmarking are intertwined within a shared framework eisert2020quantum , or concepts such as metrics and benchmarks are sometimes used interchangeably SoKBenchmarkingQuantumComputer . In Ref. SoKBenchmarkingQuantumComputer , the term physical benchmarkings has been introduced to describe properties of a quantum processor that are more accurately understood as metrics. Additionally, the concept of aggregated benchmarks has been presented, which combines metrics and benchmarking techniques. For instance, Circuit Layer Operations per Second https://doi.org/10.48550/arxiv.2110.14108 is a metric while Mirroring Circuits Proctor2021 ) is a technique employed in benchmarks. Other proposals more aligned with the intent of distinguishing between different terms have been presented, although their discussion has been limited. An excellent example of such studies is the article by Resch and Karpuzcu on the impact of quantum noise https://doi.org/10.48550/arxiv.1912.00546 .
On the other hand, in Ref. https://doi.org/10.48550/arxiv.2110.14108 the researchers from the IBM T. J. Watson Research Center took a different approach. They introduced the benchmarking Pyramid which classified in three levels the different benchmarks: Holistic (measuring the systems as a whole), Subsystem (measuring different aspects of the subcomponents used to assemble the system, e.g. classical control logic system), and Device (e.g. quantum gates or qubit registers). Apart from classifying them into these three level groups, they also distinguished two main factors that can be measured for each level: Quality (a property that determines the size of quantum circuits that can be faithfully executed), and Speed (a property related to the number of circuits that the quantum computer system can execute per unit of time). The hierarchization process carried out is interesting, and it accounts for a division of different types of, what we call, metrics. Still, during their work they focused on a procedure to generate a particular test for Quantum Volume, later defining metrics such as CLOPS on top of it but limiting the scope of the benchmark. Indeed, there are well-known extensions to the Quantum Volume to build a family of volumetric benchmarks that generalise these types of tests Blume_Kohout_2020 . Other remarkable attempts to build thorough benchmarks can be found in the literature, such as the Application-Oriented Performance Benchmarks from the Quantum Economic Development Consortium (QED-C)quantum_consortium ; lubinski2023applicationoriented , the proposal from Daniel et al. Mills2021 where they proposed measuring the performance based on three types of circuits (deep, shallow, and square classes) with three metrics that can be computed with classical computers (heavy output generation probability, cross-entropy difference, and $\ell_{1}$-Norm Distance), or the QUARK framework Finzgar_2022 .
There has been a lack of a standardized framework for assessing the efficacy of benchmarking within the field. The absence of a common and accepted standard to define benchmarking techniques and processes for quantum computing has given rise to a multitude of divergent terminologies, as well as inconsistent approaches in the field. This work aims to fill this gap by introducing some of the most prominent techniques, metrics, and benchmarks available as of the current writing, conducting a comprehensive comparison of these elements using the framework presented in this paper. Next, we will revisit some relevant techniques, metrics, and benchmarks, mapping the concepts to the common framework introduced in this work.
III.3.1 Techniques
Let us introduce some relevant techniques that are employed in benchmarks and metrics, but they are not benchmarks or metrics by themselves.
Mirror circuits: The concept of mirror circuits Proctor2021 ; PhysRevLett.129.150502 , inspired by Loschmidt echo, involves creating an alternative efficient classically simulatable circuit that mirrors a given quantum circuit $C$. This mirror circuit consists of preparing qubits in randomized Pauli eigenstates, running the base circuit $C$, applying a randomly selected Pauli layer $Q$, executing a reversed circuit $C^{-1}$ adjusted to implement a specific Pauli operation, and finally measuring qubits in their prepared basis. This approach is particularly effective for circuits containing Clifford and $Z(\theta)$ gates, a universal gate set widely used in quantum processors. The focus on Clifford gates –with $\theta$ being an integer multiple of $\pi/4$ in $Z(\theta)$ gates– is due to sufficiency in constructing various circuit types and the assumption of error-free implementation in the processors tested (verified processor), assuming error-free $Z(\theta)$ gates due to in-software phase updates. Through this method, each circuit in the set of mirror circuits has a straightforward and computable correct outcome, allowing the observed probabilities from these circuits to accurately represent the base circuit’s performance.
This technique can be employed in benchmarking to define a metric by generating a mirror circuit from one or more quantum circuits, probing different aspects of processor capability. The choice of base circuits, the number of qubits, the depth, and the structure, influence the benchmarks’ focus. In experiments, benchmarks were created from both disordered and highly structured periodic base circuits using native gates, respecting processor connectivity and implementing elements of the Clifford group.
Classical shadow tomography: Classical shadow tomography addresses challenges in efficient information extraction from quantum states doi:10.1137/18M120275X ; article12 ; zhou2023efficient . Traditional methods demand an exponential number of measurements for a complete description of $N$-qubit states, which is impractical as discussed in Section III.1. However, classical shadow tomography provides an efficient alternative for extracting information with a limited number of measurements when detailed knowledge of the entire density matrix is unnecessary. This technique uses randomized measurements zhou2023efficient , employing independent random unitary operations, $U$, selected from an ensemble. Before measurement, this operation is applied, yielding a classical snapshot $\sigma_{U,b}=U^{\dagger}|b\rangle\langle b|U$ with a probability $p_{U,b}$. The average snapshot $\sigma$ relates to the original density matrix $\rho$ through a quantum channel $M$ as $\sigma=\sum_{b}\mathbb{E}_{\mathcal{E}_{u}}[\sigma_{U,b}p_{U,b}]=M(\rho)$. Physical observables, denoted as $\langle O\rangle$, can be expressed as $\langle O\rangle=\text{tr}\left(OM^{-1}(\sigma)\right)$. Consequently, efficient classical computation of $M^{-1}$ and a low sampling complexity are key criteria for effective classical shadow tomography schemes. Different schemes correspond to diverse choices of the ensemble $\mathcal{E}_{U}$. For instance, utilizing Haar random unitaries and tensor products of single-qubit Haar random unitaries, one can prove that while the shadow norm $\|O\|^{2}_{\text{sh}}$ scales as $2^{N}$ for general $N$-qubit systems, it scales as $3^{k}$ for Pauli operators with a size of $k$, showing a better performance for few-body operators.
Efficient classically simulatable quantum circuits: Several families of quantum circuits can be efficiently simulated in classical computers. Particularly relevant are Clifford circuits, a specific class of quantum circuits composed entirely of Clifford gates. These gates include the Hadamard gate ($H$), the Phase gate ($S=\sqrt{Z}$), and the Controlled-NOT gate (CNOT). The efficient simulatability of these circuits is ensured by Gottesman-Knill theorem, which states that any quantum circuit comprising an initial state in the computational basis, Clifford gates, and measurement in the computational basis, can be efficiently simulated in a classical computer 10.5555/863284 ; nielsen00 .
Randomized Benchmarking (RB): Randomized Benchmarking (RB) is a technique used for characterizing the error rate of quantum processors. This technique involves applying a sequence of random gates, which can be sampled uniformly at random from the Haar distribution Emerson_2005 or uniformly at random from any unitary 2-design Dankert_2009 . RB has been employed as a technique for measuring errors, such as the average gate fidelity of Clifford gates PhysRevLett.106.180504 , the coherent or incoherent nature of the noise affecting Clifford gates Wallman_2015 , or the noise of a particular gate Magesan_2012 . While this technique is robust against state preparation and measurement noise, it only scales efficiently with the number of qubits when using Clifford gates.
III.3.2 Metrics
Let us introduce some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer. We will categorize these metrics into two types: Static and Non-static metrics.
Static metrics refer to internal features of the quantum processor that define specific performance-related properties. Once the processor is manufactured, these metrics cannot be altered. Next, we introduce the main static metrics:
Number of qubits: One of the most commonly used metrics in quantum performance evaluation is the number of (working connected) qubits, which directly impacts the quantum information processing capacity https://doi.org/10.48550/arxiv.2110.14108 . As the number of qubits increases, the potential of the quantum computer to address larger and more complex problems grows exponentially. We emphasize the importance of considering not only the quantity but also the quality of qubits, which is especially relevant in the NISQ era. Indeed, scaling quantum computers to a large number of qubits is a challenging engineering task that involves addressing various technical issues depending on the characteristics of the physical platform in use.
Qubits connectivity: Qubit connectivity is an important metric to consider because it determines the direct interactions between the qubits Holmes_2020 . Quantum algorithms require the interaction of multiple qubits through quantum gates. When qubits are not directly connected additional quantum gates could be necessary to facilitate their interaction increasing the runtime of the algorithm, circuit depth, and the risk of errors in the NISQ era. Therefore, qubit connectivity directly affects the performance of the quantum computer. Currently, physical architectures provide varying qubit connectivity options, such as nearest-neighbor connectivity Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 or fully connected connectivity QCion ; Zhang_2017 . One could provide the maximal, minimal and average connectivity, for instance. Other method of mapping the qubit connectivity into a numerical representation involves constructing a $n\times n$ matrix of normalized coupling terms for a $n$-qubit processor and compute its norm. This approach has as an advantage that it takes into account also the quality of the couplings but the norm complicates comparing processors with different sizes.
Gate fidelity: The gate fidelity refers to the accuracy of quantum operations SoKBenchmarkingQuantumComputer . It focuses on how well a quantum gate performs the intended operation compared to the ideal outcome. When a quantum gate is applied, there is a predetermined ideal output, but due to various factors like noise and imperfections in the quantum processor, the actual observed output might deviate from the expected one. High gate fidelity is essential for the reliability and precision of computations. If the gate fidelity is low, it indicates that the quantum gate is introducing errors, reducing the accuracy of quantum algorithms. However, although the concept is clear, its practical application in a quantum processor presents some complex problems. In quantum platforms where the entangling logic gates are two-qubit gates, such as superconducting circuits or photonics, one can apply one of these gates for each pair of connected qubits (two if the gate is asymmetric). Then one can consider metrics such as average fidelity, maximum fidelity, or minimum fidelity. However, it is complicated to compare any of these with other technologies where the gates are multi-qubit, such as trapped ions or cold atoms. In this case, moreover, the fidelity is substantially more costly to calculate.
Decoherence times: The decoherence times refer to the period during which the qubit can maintain its quantum coherence SoKBenchmarkingQuantumComputer . The interaction of qubits with the environment results in the loss of quantum superposition and the decay from the excited state to the ground state, characterized by the times $T_{2}$ and $T_{1}$ respectively. Longer decoherence times are preferable as they might permit deeper circuits to be run, reducing the probability of errors and improving computation accuracy. Unfortunately, coherence times by themselves are not an accurate measure of the quality of the processor. Decoherence occurs due to the coupling of the quantum system to the environment. Strongly protecting a quantum system can result in reduced manipulability and a substantial increase in the time required to operate quantum gates. Consequently, if we double the decoherence times but also double the gate operation times, this has no practical consequence on the processor’s performance, as the number of operations we can carry out will not change. Other thorny question is how to integrate the different coherent times of a qubit and those of different qubits.
Gate speeds: Gate speed is a metric that measures how rapidly a quantum computer can execute a given quantum gate van_der_Schoot_2023 . Considering decoherence times, this metric allows us to determine the number of gates that can be performed before the qubits lose coherence. By increasing gate speeds, we might deploy a greater number of gates within the same time interval, leading to more applied gates during the lifetime of the qubits. Here, the argument explained for coherence times can be inverted, since fast manipulable systems could also be susceptible to decoherence. The problems of this metrics are similar or worse to gate fidelity metric.
On the other hand, non-static metrics also characterize a performance-related property of the quantum processor. Unlike static metrics, after manufacturing the quantum processor, these metrics can be altered according to their application. Next, we present the main non-static metrics:
Quantum Volume (QV): This is a metric that quantifies the largest square (equal in width and layers) quantum circuit that a quantum processor can successfully run PhysRevA.100.032328 . Each layer consists of random permutation among qubits followed by pseudo-random 2-qubit gates. After running the random quantum circuit, the measured results should be above a certain threshold to be considered successful. Consequently, higher-quality quantum gates and qubits allegedly result in a larger QV. More specifically, ${\rm QV}=2^{D}$, where $D$ is the largest square successfully run. For example, a quantum volume 256 means that a quantum processor can use a subset of 8 qubits in the processor to successfully run a circuit with 8 sequential layers. A higher QV indicates a greater capability of the quantum processor to tackle more complex tasks, demonstrating advancements in the quality of qubits and error mitigation techniques. However, one limitation of QV is that it requires knowing the probabilities of all possible outcomes, which implies a large number of measurements that scale exponentially with the number of qubits. Another limitation is that the QV protocol exclusively considers square circuits. Therefore, QV does not offer a reliable prediction of the performance of circuits that have high depth and low width, or circuits with low depth and high width. Due to the strong limitations of this metric, IBM proposed the alternative Error Per Layered Gate (EPLG) mckay2023benchmarkingquantumprocessorperformance . This metric makes use of (Clifford) randomized benchmarking and mirror circuits techniques to compute the layer fidelity LF, and then $\text{EPLG}=1-\text{LF}^{1/n_{2Q}}$, where $n_{2Q}$ is the number of two-qubit gates, tipically taking the value $N-1$ for a linear chain of qubits. The exponent tries to compensate that the fidelity is a cumbersome measurement of the similarity between two many-body quantum states since it is exponentially suppressed with the size of the processor and, in the thermodynamic limit, two states are either orthogonal or the same. This leads to limitations regarding practicality and reliability, but they will not be analyzed here.
Q-Score: Q-score is an application-focused metric which measures the maximum number of variables that a quantum processor can optimize on a standard combinatorial approximate optimization problem (for instance, a Max-Cut Problem) 9459509 . Therefore, Q-Score is remarkably an application-oriented metric, since it focuses on solving a practical task. Another advantage is the versatility of the Q-score, as it can be executed and computed on any quantum processor, although in the NISQ era highly connected processors, such as the ones in trapped-ion platforms, are favored doi:10.1126/sciadv.aau0823 . On the other hand, this versatility does facilitate the comparison of devices between different computational paradigms, such as quantum annealers and gate-based vanderschoot2023qscore . Another drawback of this metric is that it does not impose any transpilation rule. This is especially important in this approximate algorithms run in NISQ processors, since the error allowed in more drastic transpilation rules can be hidden in the noise or in the approximation. Consequently, there could be strong artificial advantages of a certain processor just due to more drastic transpilation rules, so the metric is neither repeatable nor reliable.
Circuit Layer Operations Per Second (CLOPS): This metric, proposed as a natural extension of Floating Point Operations per Second (FLOPS) in supercomputers, quantifies the number of QV-type layers that the quantum processor can run per second wack2021quality . This metric was proposed by IBM to measure the performance of a processor in terms of speed, complementing QV as a metric, which was focused on measuring quality. Remarkably, CLOPS include also the classical processing and control times in the calculation. Therefore, a larger CLOPS value indicates a higher speed of the quantum processor in executing layers of the same type used in the QV. Therefore, this metric suffers from limitations similar to those of QV regarding scalability, making it neither practical nor repeatable. Additionally, it cannot be adapted to processors making use of multi-qubit gates, such as trapped ions, so it is also not consistent. Finally, as CLOPS include the time corresponding to control, one can improve it by faster classical electronics, without actually improving the processor, so reliability might also be in question. Due to these limitations, IBM proposed an update of this metric called CLOPSh www.ibm.com which updates the definition of layer and now the possible two-qubit gates are no longer arbitrary, but they are only the gates which are native in the processor (i.e. between qubits which are directly connected). The reason is other gates must be decomposed into native ones, jeopardizing the definition of layer as one composed by parallel gates. Although this update is reasonable, it does not fix any of the problems associated to CLOPS.
Algorithmic Qubits (AQ): Algorithmic Qubits (AQ=$N$) is an application-oriented metric that quantifies the largest quantum circuit of width $N$ and depth $N^{2}$ that a quantum processor can successfully run chen2023benchmarking . In contrast to quantum volume, the AQ metric does not concentrate on random quantum circuits but instead on six algorithmic classes, including Hamiltonian simulation, phase estimation, quantum Fourier transform, amplitude estimation, variational quantum eigensolver (VQE) simulation, and Monte Carlo sampling. The success of the quantum circuit is measured by evaluating the fidelity, which involves comparing the ideal output probability distribution to the measured output distribution on the quantum computer. It is considered successful only if it surpasses a certain threshold. Consequently, one must know a priori the outcome of the running algorithms, which is problematic for sufficiently large processors which cannot be classically simulated. Additionally, as the success is measured by the fidelity, it also shows scalability problems. As in the case of the Q-Score, no transpilation rules are imposed, which seriously affects reliability. As a positive point, this metric can be employed in different paradigms and platforms.
Cross-entropy difference and cross-entropy benchmarking (XEB): This metric has been used to quantify the correspondence between the experimental and the ideal output probability distributions of a pseudo-random quantum circuit Qsupremacy , and it has been used to calculate the XEB fidelity Arute2019 . The cross-entropy difference $\Delta H(p_{\mathrm{A}})$ is defined as
$\Delta H(p_{\mathrm{exp}})=H(p_{\mathrm{cl}},p)-H(p_{\mathrm{exp}},p),$
where $H(p_{\mathrm{1}},p_{\mathrm{2}})=-\sum^{N}_{i=1}p_{1}(x_{i})\log{p_{2}(x_{i})}$ ) is the cross-entropy between the probabilities distributions $p_{1}$ and $p_{2}$, and $N$ represents the total number of possible outputs. Additionally, the experimental, ideal, and uniform probability distributions are denoted by $p_{\mathrm{exp}}$, $p$, and $p_{\mathrm{cl}}$, respectively. In practice, we will have an experimental implementation of the Haar random unitary $U$ as $A_{\mathrm{exp(U)}}$ with an associated probability distribution $p_{\mathrm{exp}}$ and sample $S_{\mathrm{exp}}=\{x_{1}^{\mathrm{exp}},\dots x_{m}^{\mathrm{exp}}\}$. In this case, the (ideal) experimental cross-entropy difference is defined as
$\alpha\equiv\mathbb{E}_{\mathrm{U}}(\Delta H(p_{\mathrm{exp}})),$
where $\mathbb{E}_{\mathrm{U}}$ is the expected value over the Haar random unitaries distribution. This $\alpha$ is now estimated from an experimental sample of bitstrings $S_{\mathrm{exp}}$ obtained by measuring the output of $A_{\mathrm{exp(U)}}$ after $m$ realizations of the circuit, as
$\alpha\simeq\tilde{\alpha}=H(p_{\mathrm{cl}},p_{\mathrm{U}})-\frac{1}{m}\sum_{j=1}^{m}\log\frac{1}{p_{U}\left(x_{j}^{\mathrm{exp}}\right)},$
with an error of $k/\sqrt{m}$ and $k\simeq 1$. Here, $p_{\mathrm{U}}\left(x_{j}^{\mathrm{exp}}\right)$ is the ideal probability of the $x_{j}^{\mathrm{exp}}$ bitstring computed classically. With $m\sim 10^{3}-10^{6}$ number of samples, a sufficiently accurate approximation can be achieved. The (ideal) experimental cross-entropy difference yields a value ranging from 00 to $1$, which approximates the fidelity of the circuit. When the experimental output probability distribution aligns perfectly with a uniform distribution, it represents the worst-case scenario, obtaining an experimental cross entropy difference equal to zero. On the other hand, when the experimental output distribution perfectly matches the ideal distribution, the experimental cross-entropy difference is one. It is worth emphasizing that experimentally generating random circuits poses a significant challenge. Moreover, obtaining the ideal probability distribution requires classical computational resources that scale exponentially with the number of qubits, making this metric not practical. To discuss whether this metric is repeatable and reliable, it is important to distinguish between the theoretical method used to calculate $\alpha$ and the experimental implementation used to compute $\tilde{\alpha}$. In the theoretical method, the metric satisfies the repeatable and reliable characteristics. In the experimental implementation, there are two options: one can take a fixed finite set of Haar random unitaries, using always this set, or generate a new set of random unitaries every time. The first case satisfies both properties, but the second does not. However, this is more associated with the benchmarking methodology than the metric itself, so we will analyze the metric according to the theoretical method.
Hellinger distance: The Hellinger distance is a metric used to measure the similarity between two probability distributions. It has been used to quantify the capability of a quantum computer to accurately replicate the outcomes of a specified quantum circuit e24020244 . The Hellinger distance between the ideal probability distribution $P^{\mathrm{ideal}}$ and the experimental output distribution $P^{\mathrm{exp}}$ is defined by
$d(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sqrt{1-BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})},$
with the Bhattacharyya coefficient $BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})\in[0,1]$ defined as:
$BC(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}\sqrt{p^{\mathrm{ideal}}_{i}p^{\mathrm{exp}}_{i}}.$
The ideal and experimental probabilities to observe the $i$-th outcome are denoted by $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$, respectively. The number of possible outcomes is given by $N=2^{n}$, where $n$ is the number of qubits. The Hellinger distance converges to 00 for identical distributions and $1$ in cases, where there is no overlap between distributions. Obtaining the ideal output distribution involves simulating the quantum circuit on a classical computer, which makes this metric not practical. The theoretical method to compute this metric ensures its repeatability and reliability, although its practical application could present challenges.
Heavy Output Generation (HOG): The HOG is a metric that quantifies the success of a pseudo-random quantum circuit executed on a quantum computer 10.5555/3135595.3135617 . If we consider a pseudo-random circuit $U$ taken from a uniform Haar distribution, the ideal and experimental output probability of each bitstring after executing $U$ is denoted by $P^{\mathrm{ideal}}$ and $P^{\mathrm{exp}}$. A bitstring $z$ is classified as heavy in $U$, if its probability $P^{\mathrm{exp}}(z)$ is greater than the median of the distribution $P^{\mathrm{ideal}}$. The probability that a sample drawn from a distribution $P^{\mathrm{exp}}$ is heavy, named as the heavy output generation probability of $P^{\mathrm{exp}}$, is given by
$HOG(P^{\mathrm{exp}},P^{\mathrm{ideal}})=\sum_{i=1}^{N}p^{\mathrm{exp}}_{i}\delta_{i},$
where $N=2^{n}$, $n$ is the number of qubits, and $\delta_{i}=1$ if $i$ is heavy in $U$, and 00 otherwise. If $P^{\mathrm{exp}}$ approximates the uniform distribution, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx 1/2$; and if $P^{\mathrm{exp}}$ is close to $P^{\mathrm{ideal}}$, $\mathrm{HOG}(P^{\mathrm{exp}},P^{\mathrm{ideal}})\approx(1+\log(2))/2$. The $\mathrm{HOG}$ metric has been employed to develop a test that assesses the performance of a quantum computer as good if it exceeds $2/3$. One limitation of $\mathrm{HOG}$ is that it requires calculating the ideal probability classically, which makes this metric not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used for the cross-entropy difference and XEB metric, which means that it will satisfy both characteristics.
$\ell_{1}$-Norm Distance: The $\ell_{1}$-Norm Distance is a metric that quantifies the total difference between the output distribution of the quantum circuit produced by a quantum computer, denoted by $P^{\mathrm{exp}}$, and the probability output distribution of the ideal quantum circuit generated by a classical computer, denoted by $P^{\mathrm{exp}}$ Mills_2021 . The $\ell_{1}$- norm distance reads as
$\ell_{1}(P^{\mathrm{ideal}},P^{\mathrm{exp}})=\sum_{i=1}^{N}|p^{\mathrm{ideal}}_{i}-p^{\mathrm{exp}}_{i}|,$
where $p_{i}^{\mathrm{ideal}}$ and $p_{i}^{\mathrm{exp}}$ are the ideal and experimental probabilities to observe the $i$-th outcome and $N=2^{n}$ is the number of possible outcomes, with $n$ the number of qubits. The classical resources required to obtained the ideal probability distribution grow exponentially with the number of qubits, which makes this metric not practical. The theoretical method to compute it ensures its repeatability and reliability, although its practical application could present challenges.
Collision volume: This metric has been used to quantify the noise of a quantum computer using Haar random unitaries mari2023 . Collision volume involves executing a random circuit $U$, obtaining $m$ samples, and counting how many bitstrings have been sampled more than one time, what it is called collision. It can be proven that in an ideal scenario in which the quantum computer is not noisy, the number of collisions doubles the case in which the computer is completely noisy. Then, the collision volume is defined as
$\hat{\Delta}\equiv\frac{\hat{R}-N+N(1-e^{-m/N})}{N^{2}/(N+m)-Ne^{-m/N}},$
where $m$ is the number of samples, $N=2^{n}$ with $n$ the number of qubits, and $\hat{R}$ is the observed number of collisions. The expected value of $\hat{\Delta}$ is equal to $1$ when sampling from a typical random quantum state, and 00 when sampling from a uniform distribution. Using this metric, the authors propose a test called collision volume test, which indicates that the computer has good accuracy if $\hat{\Delta}\geq 1/2$. It is important to emphasize that the collision volume metric does not require a classical computer, but it requires around $2^{n/2+5}$ samples to work, which makes it not practical. To determine if this metric is repeatable and reliable, we follow the same criteria used in the cross-entropy difference and XEB metric, which means that it satisfy both quality attributes.
III.3.3 Benchmarks in NISQ Era
Let us now review some relevant benchmarks that have been employed to assess the performance of a quantum processor.
QPack: The QPack benchmark utilizes Quantum Approximate Optimization Algorithms (QAOA) to evaluate the effectiveness of quantum computers in solving problems such as MaxCut, dominating set, and the traveling salesman problem QPACK2021 . This benchmark assesses various aspects of a quantum computer, including runtime, output accuracy, and scalability. The performance of the quantum computer is assessed by contrasting the solution obtained through QAOA execution on the quantum computer with that obtained using a classical exact algorithm. Consequently, the size of the problem that can be handle is constrained by the computational capacity of the classical computer. The QPack benchmark has been used to compare different quantum computing platforms, such as superconducting and trapped ion devices QPACK_ .
Quantum LINPACK: Inspired by the classical LINPACK benchmark, the Quantum LINPACK benchmark assesses the performance of a quantum computer based on its ability to solve linear equations system using pseudo-dense random matrices. QuantumLINPACK . The creators of this benchmark introduce an input model known as the RAndom Circuit Block-Encoded Matrix (RACBEM), which serves as an appropriate generalization of a dense random matrix in the context of quantum computing. The success of the benchmark is measured by comparing the output vector to the ideal solution, incurring an exponential classical cost.
QASMBench: The QASMBench is a benchmark that includes a set of quantum circuits and metrics for characterizing circuit properties both before and after execution on a quantum processor li2022qasmbench . The set of quantum circuits are frequently used in various fields, such as chemistry, simulation, linear algebra, search algorithms, optimization, arithmetic, machine learning, fault tolerance, cryptography, and more QuantumAlgorithms . QASMBench is divided into three categories: small-scale, medium-scale, and large-scale, each defined by the number of qubits used. As an illustration, in the small-scale scenario, the performance of quantum processes is assessed by executing quantum algorithms using 1-5 qubits such as Quantum Walk, Linear Equation, Search, and optimization, among others. In the medium-scale and large-scale scenarios, the performance of quantum processes is evaluated by running quantum algorithms using 6-15 and more than 15 qubits, respectively. The authors propose a set of circuit evaluation metrics: Circuit Width, Circuit Depth, Gate Density, Retention lifespan, Measurement Density, and Entanglement Variance, which are indicators that allow us to quantify the performance of a certain quantum circuit executed on a quantum processor. QASMBench has been assessed on IBM-Q machines, and the effectiveness of the quantum circuits is determined by computing the Hellinger distance.
SupermarQ Suite: SupermarQ Suite is a benchmark that includes a set of applications that assess the performance of a quantum computer based on its ability to generate entanglement between qubits through the GHZ algorithm, demonstrate the quantum nature of the system through the Mermin-Bell test, solve combinatorial optimization problems through the quantum approximate optimization algorithm, find the ground state energy of the one-Dimensional Transverse Field Ising Model (1D TFIM) using the Variational Quantum Eigensolver algorithm (VQE) and simulate the time evolution of the 1D TFIM Hamiltonian using a quantum circuit via Trotterization for a specific number of time steps 9773202 . The performance of each application is assessed through different scoring function, comparing experimental outcomes with predefined ideal results. As these ideal results are known in advance, the scalability of this benchmark is ensured. The SupermarQ Suite has been employed to assess the performance of both superconducting and trapped ion devices.
QED-C Benchmark: The Quantum Economic Development Consortium (QED-C) introduced the QED-C benchmark to assess the efficiency of quantum computers lubinski2023applicationoriented . This benchmark comprises various algorithms, including Deutsch-Josza, Grover’s Search, VQE, Phase Estimation, Bernstein-Vazirani, Quantum Fourier Transform, and other algorithms. Unlike the Quantum Volume metric, which only uses square circuits, the QED-C benchmark incorporates circuits that vary in both depth and width. It includes circuits with high depth and low width, as well as circuits with low depth and high width. The QED-C benchmark evaluates computer performance by assessing fidelity, which involves comparing the ideal probability distribution with the experimental probability distribution. Therefore, this approach is limited by its requirement for exponentially costly classical computations to evaluate the performance of quantum computers.
Chemistry simulation: The Chemistry Simulation benchmark is designed to assess the performance of quantum computers by the electronic structure calculation of molecules using variational methods QuantumChemistry . The variational quantum eigensolver (VQE) algorithm has been used to calculate electronic structures using various quantum computing architectures both photonic VQE-Photonic , superconducting circuits PhysRevX.6.031007 and trapped ions PhysRevX.8.031022 . This benchmark assesses the performance of quantum computers by using ground state energy as a metric, comparing results obtained on classical computers with those derived using the VQE algorithm on quantum computers.
We have introduced the main proposals of metrics and benchmarks in the literature. Table 1 analyze metrics, evaluating which quality attributes they fulfill according to the framework introduced in Sec. II. We emphasize that we evaluate the quality attributes fulfilled by metrics in the NISQ era; Therefore, these evaluations are not valid for the post-NISQ era. We hope that as quantum processors evolve, benchmarks will adapt to each era.
Quantum Computer Benchmarking Metrics Review
Metric Practical Repeatable Reliable Linear Consistent Performance
Number of (Working Connected) Qubits Scalability
(Average) Processor Connectivity $\times$ $\times$ Scalability / Speed
(Average) Gate Fidelity $\times$ -- $\times$ Quality
(Average) Decoherence Times $\times$ -- $\times$ Quality
(Average) Gate Speed $\times$ $\times$ Speed
Quantum Volume $\times$ $\times$ $\times$ $\times$ Scalability / Quality
Q-Score $\times$ $\times$ $\times$ Scalability / Quality
CLOPS $\times$ $\times$ $\times$ $\times$ Speed
Algorithmic Qubits $\times$ $\times$ $\times$ Scalability / Quality
Cross Entropy Benchmarking (XEB) $\times$ -- Quality
Hellinger distance $\times$ -- Quality
Heavy Output Generation $\times$ -- Quality
$\ell_{1}$-Norm Distance $\times$ -- Quality
Collision volume $\times$ -- Quality
Table 1: Summary of quantum computing metrics in NISQ era. In this table, we show some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer, as presented in subsection III.3.2. We assess whether these metrics satisfy the quality attribute proposed in the literature for an ideal performance metric, as discussed in subsection II.3, although some of them are borderline. The check mark indicates that the metric satisfies the attributes; otherwise, we use the cross mark. The dash mark indicates that it is complicated to define if the metric satisfies the criterion because we cannot characterize it formally. Performance definitions: Scalability refers to the size of the processor; Quality evaluates how close the quantum processor is to a fault-tolerant quantum processor; Speed concerns the time required to process a certain task.
IV Roadmap for Quantum Benchmarking
Building upon earlier sections, our focus shifts towards a more practical exploration of guidelines when designing quantum benchmarks. Our goal in this section is to assist practitioners in evaluating existing benchmarks, making informed choices, and uncovering potential blind spots. To achieve this, we compile practical recommendations and guidelines to steer practitioners in assessing the effectiveness of current benchmarks or when considering the development of new ones. Additionally, we provide a comprehensive view of the benchmarking reporting structure, enabling researchers to convey the true essence of their benchmarks effectively. Finally, we propose the creation of an organization focused on Standard Performance Evaluation for Quantum Computers (SPEQC).
IV.1 Guidelines for the benchmark practitioner
Here, we present a set of high-level guidelines to assist benchmark designers in making informed decisions when creating benchmarks that accurately reflect the performance of the assessed quantum devices. These guidelines build upon existing industry standards and take into account insights from classical benchmarking sources (as elaborated in Sec. II), being system-independent:
Guideline 1. Benchmarks must uphold the key quality attributes: relevance, reproducibility, fairness, verifiability, and usability. In Subsec. II.1, we introduced a set of quality attributes initially suggested as guidelines for classical benchmark designers. We aim to extend the application of these attributes to the field of quantum computing devices. Relevance aids in defining the specific area within the total behavior space targeted by our benchmark proposal and clarifies how accurately it captures this domain. Fairness is key for ensuring impartial comparisons across diverse gate-sets and processor architectures. Reproducibility and usability guarantee that results obtained by a particular device running the benchmark can be verified by other vendors, fostering trust among stakeholders. Finally, verifiability is instrumental in confirming the correctness of the results. A benchmark should be self-verifiable, to enable other parties to repeat the executions and obtain similar results.
Guideline 2. Benchmarks must offer peak and base performance results. The classical baseline debate with arguments against and for an extended use of optimization flags can be extended to quantum devices. We encourage benchmarks practitioners to follow a similar approach as the final compromise established for SPEC in January 1994: ”Every SPEC CPU benchmark measurement has to measure performance with a restricted baseline set of flags, and optionally with a more extended peak set of flags”SPECHPGBenchmarksHistory . Having both metrics, peak and base, allows different vendors to share a common baseline for comparison that represents also the average non expert consumer, while enabling also the publication of the peak performance with optimised results leveraging expert knowledge over the architectures.
Guideline 3. Different technological eras necessitate different types of benchmarks. In the NISQ era, where there is no dominant architecture for quantum devices, specification-based benchmarks are particularly relevant. Given the substantial impact of noise, focusing on synthetic, microbenchmarks or kernel benchmarks could offer more practical and useful insights. As technology involves, Partially Corrected and Fault Tolerant quantum computers may transition into hybrid or kit-based benchmarks. Ultimately, for Fault Tolerant and Fully Functional Quantum Computers eras, application-based benchmarks would emerge as a reasonable and effective choice. As quantum computers progress, benchmarks will evolve to reflect this.
Guideline 4. Benchmarks metrics must uphold the key quality attributes: practical, repeatable, reliable, and consistent. They should also preferably be independent and linear. As discussed in Sec. II, for a metric to be useful in build a good benchmark it must present attributes that correlate later with quality benchmark attributes. Additionally, while linearity and independence are desirable metric attributes, they are not strictly required for effectively capturing the performance of a quantum device.
Guideline 5. Do not rely solely on a single benchmark to measure performance. While benchmarks serve as valuable tools for modeling specific aspects of the behavior of a quantum computer, relying solely on individual benchmarks may result in a lack of depth. The risk of spreading benchmarks too thin clearly arises when attempting to cover the entirety of the behavior of a device space with a single metric. Different behaviors often require distinct testing and experimentation setups, and consequently, an approach based on benchmark suites is recommended. Benchmark suites provide a more comprehensive assessment, enabling practitioners to really grasp how the computers perform in different situations.
IV.2 Benchmarking structure
Beyond adhering to the guidelines outlined in the previous section, we present a high-level step-wise benchmarking structure that encompasses the primary aspects benchmarks practitioners must consider. We anticipate that this general schema will promote a standardized approach to benchmarking, facilitating streamlined comparison efforts among various alternatives in the future, while also ensuring consistency in meeting a shared quality standard. This step-wise schema is detailed in Fig. 2.
Device verification is a crucial process to ensure that the device is working properly. Just as vendors in classical computing certify the validity of their products before consumer release, similar steps must be taken in the quantum computing field. However, the techniques for verification differ due to the fundamentally different nature of quantum computers, rooted in quantum mechanics. Since we are in NISQ era, the concepts of verification and benchmarking overlap because we can not ensure that a quantum computer will work properly. An example of device verification method is the cross-entropy benchmarking (XEB), which has been employed to verify quantum processors based on superconducting circuits Neill_2018 ; Arute2019 .
After the device verification, we proceed with the benchmark definition. First, we define the benchmark or benchmark suite that we will execute to measure the performance of the quantum processor for a certain task and define static or non-static metrics that capture the behavior of interest of the device. To execute the benchmark, users select a compiler to transform it into a quantum circuit, optimizing and converting it into an executable version on the quantum processor, considering constraints such as native supported quantum operations and qubits connectivity. Additionally, compilation rules to be allowed must be defined for both base setup and peak setup to ensure we have a baseline to compare benchmark results, as well as allowing the comparisons for peak performance when custom or expert modifications are applied. Finally, it must be established which verification protocol will be used to confirm that the benchmark has been correctly executed.
Once the benchmark has been defined, we continue with its execution. We execute the benchmark suite defined using the base and setup configurations, and measure the metrics defined for the benchmark. It is advisable to run the benchmarks multiple times and take the average of the results to reduce the impact of errors. Sometimes, the benchmark suite may not execute as expected due to a variety of reasons. Thus, it is important to verify the benchmark results using the verification protocol to ensure the benchmark is executing properly.
After completion of the benchmark execution and verification, we proceed to report the results. We must elaborate a reports compiling all results from the execution of the test suites for the peak and base configuration. Additionally, it should include the static and non-static metrics, the verification protocol, the configuration used for peak and base executions and execution process, so that others users can reproduce and verify your results. This report should be clear and contain as much information as possible so that users know how the computers are performing and can identify areas where the computers are underperforming.
Figure 2: Quantum benchmarking structure. We introduce a step-wise structure to benchmarking, from pre-benchmark steps for verification of the device, defining the benchmarking and executing it, up to the reporting stage.
Note that while the proposed step-wise approach for quantum benchmarking bears resemblances to the classical benchmarking paradigm, a closer examination reveals significant distinctions. In particular, the specific tasks required to complete each step are notably distinct.
For example, in quantum computing, we are still figuring out how to define the transpilation process and what configuration/flags are available, therefore it is not yet as developed as the compilation techniques used in classical computing. Also, unlike transpilers that maintain the abstraction level, compilers inherently involve abstraction level transformations from high-level programming languages to assembly primitives. Nonetheless, both face the challenge of arranging a primitive set of instructions optimally for the device they target. Still, because of the different goals and rules involved in compilation and layout rearrangement, unfortunately, there is no direct transferability of the knowledge.
The structure outlined provides a roadmap for pinpointing areas that require further investigation to clarify the future of quantum computer benchmarking. Researchers and benchmark practitioners are actively exploring each of the tasks, aiming to enhance our understanding and methodologies in this field. Below, we outline the primary challenges and unresolved issues that must be addressed to establish a more robust and comprehensive benchmarking process for quantum computers:
  • Device and routine execution verification. Post-NISQ era, a significant challenge we must address, which is becoming increasingly apparent with advancements in available quantum computers even within NISQ, is verifying the correctness of the results they produce. As mentioned earlier in Sect. III, we know that the problem to verify if every language in BQP admits an interactive proof where the prover is in BQP too, and the verifier is in BPP is still opened AaronsonPrize . This question has been thoroughly examined within the field, even extending to inquiries about the falsifiability of quantum mechanics as a theory aharonov2012quantum . Still, it is not the only approach explored for verifiability so far in the field, with other protocols explores with single-provers and multi-provers either quantum or classical Gheorghiu2018 .
  • Definition of good metrics. In this study, we have introduced in Subsec. II.3 the quality attributes that good metrics should ideally meet. Additionally, we have examined various proposals put forth by researchers in the field in Subsec. III.3. However, we contend that further work is necessary to precisely define the relevance of the metrics. In other words, we need to determine which aspects of the devices behavior space are being captured by these measurements. Establishing a correlation between metrics and their actual relevance to performance is crucial for obtaining diverse perspectives of the device, and a thorough assessment. Throughout the NISQ era, prioritizing metrics centered on scalability, accuracy, and quality could drive the advancement of more robust and resilient devices, as speed currently does not hinder experimentation. However, speed may become a necessary performance indicator as we transition towards the Fault Tolerant era or implement additional partial quantum error correction methods.
  • Standarise transpilation and compilation process. While we have discussed in detail the problematic regarding the transpilation, which can be mapped to the idea of how to optimise the intermediate representation (IR) language descriptor for quantum programs (e.g. OpenQASM cross2017open or QASM 1580386 ), we have not detailed the mapping into the physical layer. This aspect of the compilation and transpilation funnel relates to the Quantum Physical Operations Language (QPOL) 1580386 . Due to the lack of a standard architecture for the development of the quantum technologies and the complex nature of simplifying the intermediate representation (IR) of circuits, much of the research focus has revolved around this aspect. In the future, we might have to include the exploration of QPOL optimization techniques through standard compilers within the industry. Even today, transpilation alone is witnessing a surge of initiatives aiming to delve deeper into its complexities. Techniques for gate decomposition, routing, and mapping are continuously evolving, alongside other approaches have further explored ideas such as parameterized circuit instantiation where a feedback loop is allowed 9951320 . No standards or rules are defined to what is the fair use of these techniques for quantum devices. From the classical computing benchmarking past, we know good examples of techniques that are not allowed to encourage a fair competition of benchmark results. For instance, the definition of assertion flags is not allowed, i.e. flags that assert a certain property of the program, one that may not hold for other routines. Another example of a restriction to the configuration of a classical benchmark execution is that no more than two compilation passes without processing in between are permitted SPECHPGBenchmarksHistory . We believe on the quantum computing field there is still a need to further explore the fair usage of optimization routines and flags on transpilations to standarise the comparison of benchmark results for different devices, resembling these previous efforts on classical benchmarking. This debate extends too to the process of establishing peak and base configurations.
  • Matching routines to relevant regions of the behaviour space. Even though we are still in the early ages of quantum computing, there is a huge investment to find and identify real-world applications that can benefit from this technology. This search is particularly relevant for defining which regions of the behavior space of a quantum device need to be evaluated. This effort needs to be recurrent, to prevent us from falling prey to the Goodhart’s law as warned before. We firmly believe an active effort to keep uncovering what aspects of the computational device are useful is required to keep modelling new benchmarks accordingly. These new discoveries will make some of our current benchmarks obsolete, while opening the door for future benchmarks to be created. As Hennessy and Patterson pointed out in their classic book on computing architecture, it is fallacious to assume that benchmarks retain their validity indefinitely Hennessy2011-kt .
IV.3 Standard Performance Evaluation for Quantum Computers (SPEQC)
As we have discussed in this article, assessing the performance of a quantum processor presents a challenge without a straightforward solution. This underscores the risks of relying solely on a single benchmark or, even worse, relying on a single metric. Therefore, it is essential for us to collectively establish a standardized set of good benchmarks. These benchmarks would facilitate effective evaluation and comparison of different quantum processors for both users and manufacturers.
Leveraging the lessons learned and inspired by the creation of Standard Performance Evaluation Corporation (SPEC) in classical processor benchmarking, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC). Drawing inspiration from the founding principle of SPEC, ”let us no longer use small, synthetic programs as benchmarks, let us use real application programs” SPECHPGBenchmarksHistory , should aim to define metrics similar to the SPECmark. The proposed metrics should avoid replicating what was done with MIPS (Millions of Instructions Per Second) in the 1980s. At that time, MIPS was recognized as an inaccurate measure of actual computer performance, leading to the development of more appropriate metrics. Similarly, SPEQC’s metrics should aim to replace current efforts to establish a standard metric to measure the performance of quantum computers.
SPEQC should be a non-profit organization dedicated to establishing criteria for evaluating the performance of quantum computers, as well as developing and distributing a standardized set of tests to be used as benchmarks. Furthermore, SPEQC should provide the industry with clear specifications to assist benchmark designers in creating assessments that accurately reflect the performance of quantum devices. In general, these specifications should follow the guidelines for the benchmarks introduced in Subsec. IV.1.
Performance evaluation standards provide a common framework for assessing the performance of quantum computers, reducing the potential for bias in the evaluation process. When evaluation criteria are clearly defined and applied uniformly, it becomes easier to compare different results and objectively determine the performance level of the devices. Additionally, performance evaluation standards help maintain consistency in the evaluation process, ensuring that all quantum devices are evaluated using the same criteria. This consistency is essential for guaranteeing fairness and avoiding any perception of favoritism towards a particular architecture.
Establishing standards for execution, compilation, and transpilation, as well as rules for reporting results and standards, would ensure consistency, clarity, credibility, and, most importantly, the ability to reproduce performance results. Fig. 3 shows an example of how quantum processor performance results could be reported. The report would offer detailed information about the hardware and software, including the processor name, native gates, transpiler, and other relevant details. Furthermore, the results tables provide the performance of the quantum processor when executing different benchmarks. We emphasize that this results report is an example and may contain more information. These reports should be accessible to provide consumers with transparent information about manufacturers product performance, facilitating informed decision-making across the industry.
The development of benchmarks is a complex task and requires making significant decisions that must satisfy several design objectives simultaneously. In other words, when designing benchmarks, it is necessary to balance different objectives that are important for the design and functionality of the benchmark. These objectives may include easy to use, relevance, and fairness, among others. One of the most relevant aspects to highlight is the need to ensure that the benchmarks developed are fair. To achieve this, SPEQC should actively involve and collaborate with a diverse range of stakeholders. This would involve participation from representatives various sectors, including the scientific community, industry, public administration, and hardware and software vendors. A benchmark developed with the participation of all stakeholders is considered fairer than one designed exclusively by a single company. Although involving all stakeholders may not be the most efficient method for developing a benchmark, consensus among multiple stakeholders will make the benchmark fair. Consequently, SPEQC should adopt an inclusive membership policy, whereby any company or organization with an interest in advancing the field of quantum computing is encouraged and welcomed to join the initiative. This inclusive approach would ensure that the development and implementation of benchmarks are reflective of the broader interests and needs of the quantum computing community, thereby maximizing their impact and relevance in the industry. By promoting an open and collaborative environment for debate, the exchange of knowledge, perspectives, and best practices is facilitated, driving innovation and progress in the field of quantum computing.
Figure 3: Example of SPEQC report. Report inspired by the SPEC CPU® 2017, none of the metrics are reflecting actual measurements of the devices or real properties, their usage is limited to exemplify how the report would look. More details can be added to the report file in the form of sections below Submit Notes.
V Conclusion
In summary, we have proposed guidelines to assist benchmark developers in creating fair and good benchmarks. Drawing from extensive experience in the field of classic benchmarks, we have defined the quality attributes that must be met for a quantum benchmark to be considered good. It is important to note that meeting all these attributes simultaneously can be challenging. We recommend using benchmark suites that include different tests to assess the performance of quantum computers in various tasks. These suites offer a more comprehensive evaluation, allowing professionals and users to truly understand how quantum computers operate in different work scenarios.
Additional, drawing from past lessons, we critically reviewed the existing methods and strategies proposed for quantum benchmarking. We identified inconsistencies in the terminology and concepts used in this field. To address this, we propose adopting a standardized vocabulary for quantum computer benchmarking. This common language will help structure our collaborative efforts more clearly, enabling more accurate measurement of quantum performance.
Along with proposing guidelines for benchmark designers, we have introduced a benchmarking structure that includes the essential elements that benchmark professionals should consider. This structure provides a clear guide on the necessary steps to ensure a reliable and fair comparative evaluation process. Identifying the essential elements helps avoid biases and establishes quality standards for evaluating quantum devices. Furthermore, this structure highlights areas that require more attention to advance the field of quantum computing benchmarking.
Standardization in the evaluation of quantum device performance is essential for fair and objective assessments, ensuring the quality and reliability of test results. To this end, we propose the creation of SPEQC, a nonprofit organization dedicated to establishing clear criteria for evaluating quantum computers. By defining these criteria, biases and subjectivities are eliminated, resulting in more accurate and reliable evaluations. Our focus on advancing the standardization of quantum device performance evaluation aims to drive progress in quantum computing technology in the right direction, while acknowledging the risks of falling prey to Goodhart’s Law. For this reason, we expect that our benchmarking framework, along with continuous efforts from the community, will help us stay on the right track, always vigilant for potential pitfalls in our approaches.
VI Acknowledgments
We thank T. Monz, A. Frisk-Kockum, Z. Zimboras, and F. Wilhelm-Mauch for their helpful discussions and valuable comments, and for carefully reading our article. The authors acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, as well as from the EU FET-Open project EPIQUS (Grant No. 899368), also from Project Grant No. PID2021-125823NA-I00 595 and Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future,” this project has also received support from the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call –Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan– NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge funding from Basque Government through Grant No. IT1470-22, through the ELKARTEK program, project KUBIT (KK-2024/00105), and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This work has been supported by Basque Government through the BERC 2022-2025 program.